# Microseismic event classification with a lightweight Fourier Neural Operator model

### v1

@@ Introduction
Figure 2: F1 score versus probability threshold results for the test set of 500 random signal and 500 random noise samples from the microseismic dataset. Six deep learning models are evaluated. Trained on a limited number of data, the FNO-based model achieved an F1 score of 0.98.
Results
FNO Model for Earthquake Detection
The training process of the earthquake detection FNO model was conducted over 8 epochs. Here, we show the results for the test set that the trained model did not see before (blind testing). Figure 3 shows the number of true positives, false positives, false negatives, and true negatives obtained by comparing the model output and the labels. The model’s performance was measured by metrics such as accuracy, precision, recall, and F1 score. The achieved accuracy on the test set was 95%, precision 94%, recall 96%, and F1 score 95%. These results demonstrate the model’s ability to capture data features even with a limited training set of only 100 waveforms. Additionally, the model achieved human-like accuracy in a much shorter time, demonstrating its robustness and efficiency.
Figure 3: Confusion matrix results for the test set of 1000 random signal and 1000 random noise samples from the STEAD dataset. Trained on a limited number of data, the FNO-based model achieved an F1 score of 0.95.
Microseismic Data Testing
@@ Results
In practical scenarios, models trained on one dataset often need to be applied to different, previously unseen datasets. Performance in such cross-domain applications may differ significantly from performance observed within the original training domain. These differences arise both from variations in seismic data characteristics between the training and target datasets and potential biases introduced by specific annotation methods or data selection processes during training. To assess model robustness in cross-domain conditions, we conducted a cross-evaluation by testing each pre-trained model on the microseismic test dataset (Figure 4). Choosing the optimal threshold, our FNO-based model and PhaseNet achieved the highest performance (F1 = 0.98), closely followed by EQTransformer (F1 = 0.95) (Figure 2). Models with narrower input windows, such as DPP (F1 = 0.85), CRED (F1 = 0.79), and GPD (F1 = 0.67), showed significantly lower scores. This large performance gap highlights the advantage of models such as PhaseNet and EQTransformer, which analyze longer waveforms and thus benefit from a broader temporal context, such as moveout and reflections, to improve phase-identification accuracy.
Figure 4: Confusion matrix results for the test set of 500 random signal and 500 random noise samples from the microseismic dataset. Six deep learning models are evaluated. Trained on a limited number of data, the FNO-based model achieved an F1 score of 0.98.
@@
EQTransformer performs better than CRED, which can be explained by the enhanced architecture: EQTransformer is an extended version of CRED that uses attention structures in addition to the CNN and RNN. The poorer performance of DPP may be attributed to its hyperparameters being specifically optimized for a local seismic dataset exhibiting significantly different seismic characteristics. Although the other evaluated models were initially developed using local or regional datasets, their hyperparameters have not undergone the same level of systematic optimization as the DPP model, which may make them more robust when applied to diverse seismic data.
@@
Discussion
An important factor when choosing deep learning models is their computational requirements. As in most applications, inference time matters more than training time; we do not perform this analysis here.
Deep neural networks such as EQTransformer and PhaseNet have significantly advanced earthquake detection and P- and S-phase picking accuracy. However, the complexity of these models limits their practical application in real-time, especially for microseismic monitoring of well operations, which requires lightweight and efficient processing. To address this limitation, we developed an FNO-based detection model, an optimized yet high-performance deep learning model suitable for resource-constrained environments. Training the FNO model typically takes several minutes, depending on the complexity of the operator model and the size of the training set. However, the inference (or testing) time takes less than one second. This makes our approach highly efficient and well-suited for real-time seismic data processing.
Conclusion
We developed and tested a lightweight model based on the Fourier Neural Operator (FNO) for seismic event detection. The model has clear benefits over established automated methods and deep-learning models that need more computational resources. We tested it initially on the STanford EArthquake Dataset (STEAD), and our model performed well (F1 = 95%) even with small training data. Cross-testing with additional microseismic datasets confirmed the robustness and effectiveness of our FNO model. It achieved an F1 score of 98%, reaching and outperforming baseline models such as PhaseNet and EQTransformer, respectively. Our FNO model is lightweight and requires less computational power, a significant advantage when processing real-time data. This makes it very suitable for use in computationally constrained environments such as microseismic monitoring. We show that network detection based on a minimum number of stations that detect a seismic event requires careful choice of the threshold for detection that minimizes the number of false detections while maximizing the number of true detections. FNO methodology appears to be very robust and easy to implement for this purpose.

### v2

@@ INTRODUCTION
Figure 2: Sensitivity analysis of the F1 score for the STEAD test set with respect to the probability threshold for the six evaluated deep learning models. The curves illustrate the dependency of detection performance on the chosen decision boundary across the full probability range [0, 1]. Star markers indicate the optimal threshold identified to maximize the F1 score for each respective architecture.
RESULTS AND DISCUSSION
STEAD Dataset Performance
To verify the FNO model’s effectiveness under data-sparse conditions, we first evaluated its performance across training set sizes ranging from 100 to 500 samples (Figure 3). The results indicate that the FNO model maintains a robust F1 score above 0.93 even with minimal training data, with performance stability significantly improving as the dataset size increases beyond 200 samples (Table 2). For the final performance evaluation and comparison with the results of the other methods, the model trained on the largest dataset (500 samples) was used. The hyperparameters of the model were selected based on the lowest validation loss: modes = 16, width = 32.
Figure 3: Impact of training dataset size on the classification performance (F1 score) of the FNO model. The box plots illustrate how F1 scores are distributed for different model setups (different modes and widths) for different training set sizes (ranging from 100 to 500 samples). The line in the middle of each box marks the median F1 score. The edges of the box indicate the interquartile range. The line with markers tracks the mean F1 scores, illustrating that the accuracy increases and stability improves (reduced variance) as the number of training samples increases.
Table 2: Performance statistics (F1 score) of the FNO model across varying training dataset sizes.
Training Size Mean F1 Std Dev Min F1 Max F1
100 0.9305 0.0093 0.8942 0.9409
200 0.9426 0.0050 0.9288 0.9493
300 0.9434 0.0047 0.9339 0.9501
400 0.9434 0.0053 0.9342 0.9539
500 0.9488 0.0054 0.9306 0.9565
For the testing of the model’s ability to generalize, a blind test was conducted on 1,000 independent waveform windows ($N_{noise}=500$, $N_{signal}=500$). Table 3 shows the performance metrics for the FNO model and the five different architectures. The FNO-based model’s accuracy was 95.2%, and the F1 score was 0.953. The highest metrics were for the EQTransformer and the PhaseNet models, with F1 scores of 0.999 and 0.960, respectively. It is important to note that these baseline models benefit from extensive pre-training on large-scale global datasets. However, the FNO model shows a strong ability to detect the main features of the waveforms and deliver high-quality results with a limited training set of only 500 samples.
The error distribution in Figure 4 further shows the performance of the different models. The FNO architecture maintains a balanced performance with 31 false positives and 17 false negatives. This suggests that the FNO architecture is not biased towards either class. In contrast, the CRED and DPP models show a high number of errors. These results confirm that the FNO model is a robust and efficient alternative for seismic event detection, achieving high human-like accuracy and generalization ability comparable to established benchmarks.
Table 3: Classification performance metrics for the six deep learning models on the test set of 1,000 samples (500 noise, 500 signal) from the STEAD data. The best performing model in each category is highlighted in bold.
Model Accuracy Precision Recall F1 Score False Positives False Negatives
EQTransformer 0.999 0.998 1.000 0.999 1 0
PhaseNet 0.960 0.956 0.964 0.960 22 18
FNO 0.952 0.940 0.966 0.953 31 17
CRED 0.785 0.748 0.860 0.800 145 70
DPP 0.721 0.651 0.952 0.773 255 24
GPD 0.506 0.503 0.988 0.667 488 6
Figure 4: Error distribution analysis of the six DL models on the STEAD test set. The bar chart shows the False Positives (Type I errors, red) and False Negatives (Type II errors, blue). While PhaseNet and FNO demonstrate balanced and minimal error rates, CRED shows a high tendency for false alarms, whereas DPP is prone to missing valid seismic signals.
Microseismic Data Performance
@@
In practical scenarios, models trained on one dataset often need to be applied to different, previously unseen datasets. Performance in such cross-domain applications may differ significantly from performance observed within the original training domain. These differences arise from variations in seismic data characteristics between the training and target sets and possibly some bias in the annotation process and the selection of the training data. To validate the performance of the model in different domains, a cross-evaluation was performed by testing each of the pretrained models on the microseismic test set (Table 4, Figure 5).
Table 4: Classification performance metrics for the six deep learning models on the blind test set of 1,000 samples (500 noise, 500 signal) from the microseismic data. The best performing model in each category is highlighted in bold.
Model Accuracy Precision Recall F1 Score False Positives False Negatives
FNO 0.980 0.992 0.968 0.980 4 16
PhaseNet 0.982 0.994 0.970 0.982 3 15
EQTransformer 0.946 0.963 0.928 0.945 18 36
CRED 0.761 0.700 0.914 0.793 196 43
DPP 0.862 0.915 0.798 0.853 37 101
GPD 0.500 0.500 1.000 0.667 500 0
Figure 5: Error distribution analysis using microseismic test data for six models. The bar chart represents False Positives (Type I errors - red) and False Negatives (Type II errors - blue).
Choosing the optimal probability threshold, our FNO-based model and PhaseNet achieved the highest scores (F1 = 0.98). EQTransformer was very close with an F1 of 0.95. Models with smaller input window size, such as DPP with an F1 of 0.85, CRED with an F1 of 0.79, and GPD with an F1 of 0.67, scored significantly lower. This highlights that our model, PhaseNet, and EQTransformer have an advantage due to their ability to analyze longer waveforms and have a broader context over time, making them perform better in event detection.
@@
EQTransformer outperforms CRED because of the enhanced architecture of the model. EQT is an extended version of the CRED model that uses an additional attention component with CNN and RNN. DPP performs poorly because its hyperparameters were tuned for a local dataset that has very different seismic features. Although the other evaluated models were initially developed using local or regional datasets, their hyperparameters have not undergone the same level of systematic optimization as the DPP model, which may make them more robust when applied to diverse seismic data.
@@
It is important to clarify that while this constitutes a cross-domain application, the source magnitudes in our target dataset ($0.5<M<2.5$) significantly overlap with the training distributions of most baseline models (typically $0.5<M<3.5$; see Table 1). Therefore, any performance degradation observed in the benchmark models is less likely attributable to a discrepancy in event magnitude and more likely driven by differences in sensor characteristics between the regional seismic networks and the local microseismic monitoring array.
Waveform Analysis and Error Characterization
To investigate the performance disparities observed in the statistical evaluation, we conducted a qualitative analysis of the waveform characteristics and specific failure modes. Figure 6 illustrates a domain mismatch between regional tectonic earthquakes in STEAD and locally recorded induced seismicity. The datasets differ systematically in recording scale (local vs. regional), propagation path/site response, and noise conditions, which together contribute to the observed performance gap. This discrepancy is a primary driver of the generalization failures observed in Figure 7, where the baseline PhaseNet model, optimized for longer-period tectonic events, may miss clear microseismic arrivals or misidentify high-frequency noise resulting from anthropogenic activity as seismic signals.
Figure 6: Comparison of example waveforms from the two datasets, illustrating their differing time–frequency characteristics. (a) STEAD tectonic earthquake. (b) Microseismic event from this study.
Figure 7: Analysis of classification failures by the baseline PhaseNet model on the STEAD and microseismic test sets. (a) A False Positive on STEAD, where instrumental noise is misclassified as a tectonic event. (b) A False Negative on STEAD, which indicates that an earthquake was missed. (c) A False Positive on the microseismic data. Although the models have similar probability scores for this noise data, PhaseNet identifies it as an event (above its optimal threshold $\tau=0.85$). However, the FNO model rejects it, using a stricter threshold ($\tau=0.90$). (d) A False Negative on the microseismic dataset, where a low-magnitude event remains undetected. Prediction probabilities ($P$) for each model are indicated in the respective panel subtitles.
In contrast, the failures of the FNO model (Figure 8) are less attributable to domain mismatch and more to the specific properties of the spectral processing. Because FNO layers rely on global Fourier transformations, they are highly efficient at capturing wave propagation patterns but can be susceptible to strong, monochromatic noise sources (e.g., resonance). In such cases (Figure 8c), the persistent spectral energy of the noise artifact overlaps with the learned frequency modes of seismic signals, leading to false positives. Furthermore, while FNO maintains high sensitivity, its lightweight architecture, comprising significantly fewer parameters than U-Net based models, may limit its ability to disentangle extremely weak signals from complex, non-Gaussian background noise (Figure 8d).
Figure 8: Analysis of classification failures by the proposed FNO model on the STEAD and microseismic test sets. (a) A False Positive on the STEAD dataset, where a complex noise burst is incorrectly flagged as a tectonic event. (b) A False Negative on STEAD, representing a missed earthquake signal, likely due to the model’s limited capacity to resolve low-amplitude onsets within the coda of preceding events. (c) A False Positive on the microseismic dataset triggered by monochromatic resonance noise; the global spectral nature of the FNO architecture makes it sensitive to such strong periodic artifacts which mimic the spectral energy of a seismic signal. (d) A False Negative on the microseismic dataset, where a weak event ($SNR<1.5$) is obscured by high-amplitude ambient field noise.
Multi-station Detection and SNR Effect
@@
Computational Cost
Deep neural networks such as EQTransformer and PhaseNet have significantly advanced earthquake detection and have been successfully deployed in real-time regional monitoring systems (e.g., the Southern California Seismic Network). However, optimization for strictly resource-constrained environments, such as downhole acquisition tools or battery-operated remote nodes, remains a valid engineering goal.
As shown in Table 1, PhaseNet and EQTransformer utilize approximately 268,000 and 377,000 parameters, respectively. In contrast, the proposed FNO-based model requires only 34,000 parameters. This order-of-magnitude reduction in model size lowers the memory footprint and computational burden, making the FNO approach uniquely well-suited for edge-computing applications where hardware resources are severely limited, or for scenarios where models must be rapidly retrained on small, local datasets.
CONCLUSIONS
We developed a lightweight Fourier Neural Operator (FNO) model for binary seismic-event classification from three-component waveform windows. The proposed architecture provides competitive detection performance while using an order of magnitude fewer trainable parameters than widely used deep-learning baselines (34,000 versus 268,000 for PhaseNet and 377,000 for EQTransformer), reducing memory footprint and inference cost. On STEAD, the FNO achieved F1 = 0.953 on 1,000 blind test windows and showed strong data efficiency, maintaining F1 > 0.93 even when trained on only 100 labeled examples. On a field microseismic dataset, the model reached F1 = 0.980 on 1,000 blind test windows after training on 200 labeled windows, closely matching PhaseNet (F1 = 0.982) and outperforming several benchmark models in cross-domain testing.
We further show that practical reliability can be improved through array-style decision-making: requiring detections to be supported by a minimum number of stations reduces false alarms, but necessitates careful selection of the probability threshold to balance missed events against false detections. Overall, the FNO-based classifier offers an efficient and accurate option for near–real-time microseismic monitoring, particularly in settings where computational resources and labeled training data are limited.